Consider the constant rate of depreciation value of a certain car. You bought a car amounting to 1.5 million pesos. After 2 years, the value decreased to 1.2 million pesos. If the amount of the car is directly proportional to the constant rate of decay in amount and such decrease in amount continues at the same rate, then how much is the estimated value of the car after 10 years?


1
Expert's answer
2022-06-09T08:10:16-0400
"\dfrac{dy}{dt}=-ky"

"y(t)=y(0)e^{-kt}"

Given "y(0)=1500000, y(2)=1200000."

Then


"1200000=1500000e^{-k(2)}"

"e^k=\sqrt{1.25}"

"y(t)=1500000(\sqrt{1.25})^{-t}"

"y(10)=1500000(\sqrt{1.25})^{-10}"

"y(10)=491520"

491520 pesos.